# A model-based assessment of social isolation practices for COVID-19 outbreak response in residential care facilities

### v1

@@ Abstract
We apply a computational modelling approach to investigate the relative effectiveness of general isolation practices for mitigation of COVID-19 outbreaks in residential care facilities. Our study focuses on policies intended to reduce contact between residents, without regard to confirmed infection status. Despite the ubiquity of such policies, and their controversial association with adverse physical and mental health outcomes, little evidence exists evaluating their effectiveness at mitigating outbreaks. Through detailed simulations of COVID-19 outbreaks in residential care facilities, our results demonstrate that general isolation of residents provides little additional impact beyond what is achievable through isolation of confirmed cases and deployment of personal protective equipment.
I Introduction
@@
Residential aged-care facilities (RACFs, also called long-term care facilities, aged care homes, or nursing homes) have elevated risks of respiratory infection outbreaks and associated disease burden. Heightened risk arises due to importation of pathogens by visitors and staff, close and prolonged contact among residents and staff, and the demographic and health profiles of residents. These intersecting factors produce scenarios in which outbreaks are difficult to mitigate and carry disproportionate consequences in terms of medical impact. Simultaneously, the physical infrastructure, insecure workforce, high health needs of residents, and particular social requirements of RACFs impose limitations on the types of control measures that can be enacted without compromising the mental health of residents [1].
@@
Modelling efforts made early during the COVID-19 pandemic did not emphasize the confluence of risk factors present in aged-care environments that would later be acknowledged as responsible for large numbers of preventable deaths [2]. This led to a substantial and ongoing global research effort applying infectious disease modelling to outbreak detection, response, and prevention in residential care scenarios [3, 4, 5, 6, 7, 8, 9].
@@
In particular, there is a lack of modelling work that specifically assesses the effectiveness of general isolation conditions applied to resident populations, despite these being possibly the most controversial policies implemented across the sector globally. We refer specifically to policies that were intended to reduce contact between residents, without regard to confirmed infection status. Around the world, though such policies were not consistently included in government recommendations, many research articles, clinical reports, interviews, media reports, and informal testimony agree that facilities implemented general isolation measures in attempts to reduce the impact of ongoing outbreaks [12, 13, 14, 15, 16, 17].
Here we apply an agent-based model of outbreak mitigation in aged care, to investigate the effectiveness of such policies, and assess the conditions under which they may be justified. For a realistic assessment, our model includes other generic interventions consistent with the Australian Government’s recommendations released during the COVID-19 pandemic: isolation of confirmed resident cases, furlough of staff with confirmed infection, and deployment of personal protective equipment (PPE) after outbreak declaration [18].
Our results indicate that general isolation of residents does not provide substantial additional benefits beyond those achieved through screening, isolation of confirmed cases, and deployment of PPE. Our results also highlight the importance of effective case isolation, and indicate that asymptomatic screening of residents and staff may be warranted, especially if importation risk from the outside community is high.
II Summary of methods
In consultation with the Australian Government Department of Health and Aged Care, we developed an agent-based model of COVID-19 transmission in a structured population, building upon and modifying earlier modelling frameworks used to assess international arrival and quarantine pathways [19]. Key features of the model include detailed representation of:
@@ II Summary of methods
To capture the limitations of early detection, we implemented realistic screening and response strategies. These were based on those set out in the Australian Government Department of Health document “COVID-19 Outbreaks in Residential Care Facilities, Communicable Disease Network Australia, National Guidelines for the Prevention, Control and Public Health Management of COVID-19 Outbreaks in Residential Care Facilities” dated February 15th, 2022 [18]. Full details of the model implementation and assumptions are provided in the Methods.
Briefly, a facility is modelled as a static population of residents, allocated to single-occupancy rooms. For this study, we simulated a facility with 121 staff and 88 residents, which is a typical size for an Australian RACF. Staff attend a facility according to a weekly roster, with each staff member working five, three, or two days per week which accounts for a total of 73.8 full-time equivalent (FTE) staff (in Australia, 1.0 FTE corresponds to 5 full-time shifts in one week). Staff members are allocated to a set of rooms housing residents they will visit during the days they attend a facility. These allocations are used to generate a network of potential contacts, from which potentially infectious contact events are randomly sampled (Figure 1). In addition to these structured contacts, random contacts between residents in different rooms are included separately to allow for interactions mediated by communal areas and group activities that are not explicitly simulated in the facility model. Infectious contacts are subject to a probabilistic transmission process that depends on the time since the transmitting case was infected and any mitigation measures in place (see Methods).
Infection proceeds through an incubation period, followed by symptom expression and eventual recovery (see Methods for more details on the model of individual infectiousness trajectories). Each infected individual is potentially infectious throughout their entire infectious period, including the incubation period and the duration of symptoms. Infectiousness increases until approximately the time of symptom onset, and then declines until recovery. We assume that 33% of infections never express symptoms and that asymptomatic infections have the same distribution of infectiousness trajectories as those which do express symptoms.
@@
The model includes asymptomatic testing of staff and residents, as well as testing on development of symptoms. Infected residents are (imperfectly) isolated from other residents upon detection, but still have contact with staff. Infected staff are furloughed for 7 days upon detection, and have no further contact within a facility during the furlough period (Figure 4). As per the Australian Government’s guidelines (dated 15th February, 2022 [18]), an outbreak is declared once two cases have been detected in residents within a five day period or five cases have been detected among staff members over a seven day period (Figure 3). Once an outbreak has been declared, the frequency of testing of staff and residents may increase and infection control measures (e.g., use of PPE) are put in place, reducing the probabilities of transmission among staff and residents (Figure 5).
Figure 3: Schematic of the transmission process and outbreak declaration after detection of a specified number of cases within a given time window. Outbreaks are declared either from the detection of two cases in residents within 5 days, or detection of 5 cases in staff members within 7 days.
We calibrated the disease transmission model to produce a reproductive ratio of $R_{0}\approx 2.4$, to simulate a scenario consistent with SARS-CoV-2 transmission in populations with modest immune-derived protection. These conditions correspond approximately to early estimates of the effective reproduction number of the Omicron variant [22]. Further, the reproductive ratio we selected is broadly consistent with the assumptions used in previous modelling studies [6, 23, 3]. Under these conditions, we simulated the following set of policies for asymptomatic screening:
@@
Additionally, for each of the above scenarios, we vary the extent to which the resident population (regardless of case status) is socially isolated from one another following declaration of an active outbreak. Here, we present results for 0% (no general isolation), 50% (partial isolation), and 90% (stringent isolation) reductions in background contact rate. Note that residents in case isolation always have their background contact rate reduced by 90% (all scenarios).
@@
This produces a set of 10 scenarios. For each screening strategy (other than the unmitigated outbreak scenario) we perform simulations until 1000 outbreaks have been simulated (the total number of simulations required for this can vary depending on the screening strategy). This produces an outbreak size distribution that is unimodal because it does not include those simulations in which transmission dies out early on. For the unmitigated scenario (in which outbreaks are not declared), a fixed number of independent simulations are performed (10,000 instances). This produces a bimodal distribution of outbreak sizes because it includes those simulations in which transmission dies out early (Figure 6a). We report summary statistics for the second mode (large outbreaks) to facilitate comparison between unmitigated and declared outbreaks. For each scenario, we report the following summary statistics:
@@
  • time to outbreak declaration: the time between introduction of the index case ($t_{0}$) and declaration of the outbreak.
@@
The total number of infections is a generic measure of outbreak severity, but does not account for time-dependent properties of an outbreak. We report the outbreak duration as a measure of severity because costly and burdensome mitigation measures are in place throughout this period. The total number of residents placed into case isolation due to confirmed infection provides an additional measure of severity relevant both to apparent infection numbers and also to detrimental impacts of social isolation. The time to outbreak declaration and the time to first detection do not vary as functions of outbreak response, and provide estimates of the benefits of asymptomatic screening for early detection. The peak staffing deficit (measured in FTE) is a measure of the labour shortage (absenteeism) in the workforce due to furlough of staff. We chose to report the peak FTE deficit rather than the cumulative deficit because it provides a more intuitive measure of the logistical stress experienced within the facility due to personnel shortages. For example, if 1 staff member working 5 shifts in one week were furloughed, this would produce 1.0 FTE deficit. If one staff member working 2-days per week were furloughed, this would produce a 2/5 FTE deficit.
III Results
III.1 The effects of asymptomatic screening strategies
@@
III.2 The effects of general isolation conditions
@@
Our results demonstrate that, while partial and full isolation of the resident population can make a marginal difference in the absence of asymptomatic screening (Figure 8a), it has a negligible effect when combined with asymptomatic screening during outbreaks (Figure 8b). This somewhat surprising finding is due in part to the effectiveness of asymptomatic screening, targeted case isolation, and PPE use, and in part to a fundamental constraint to isolation of aged care residents: that residents must remain in contact with facility staff during outbreaks. Therefore, even if residents are prevented from interacting with one-another, their interactions with staff members can result in transmission. This, combined with the relatively substantial reductions in transmission from the use of PPE and the isolation of confirmed cases, leaves little to be achieved through general isolation of the resident population. For comparison with Figure 8, frequency distributions of cumulative cases for scenarios in which all outbreak mitigation procedures (PPE, case isolation, and asymptomatic testing) are disabled are shown in the Supporting Material Figure S3. This comparison demonstrates that reducing contact between residents can mitigate against transmission and makes a substantial difference to infection totals when other response measures are not implemented.
@@
screening strategy
general isolation level
cum. infections
outbreak duration
peak FTE deficit
cum. residents isolated
full asymptomatic screening
90%
14
11
3.4
6
full asymptomatic screening
50%
15
12
3.6
6
full asymptomatic screening
0%
17
14
3.6
7
asymptomatic screening during outbreaks
90%
29
14
6.6
12
asymptomatic screening during outbreaks
50%
31
15
6.4
13
asymptomatic screening during outbreaks
0%
33
16
6.6
15
no asymptomatic screening
90%
32.5
16
4.4
9
no asymptomatic screening
50%
37
20
4.2
11
no asymptomatic screening
0%
44.5
24
4.4
15
unmitigated outbreaks
$173^{*}$
$*{}^{*}$ Median over all unmitigated outbreaks which produced more than 80 infections (corresponds to the second mode of the distribution shown in Figure 6a).
Table 1: Summary statistics for model output over a range of mitigation scenarios. Medians values are shown for cumulative infections, outbreak duration, peak FTE deficit, and the cumulative number residents placed into case isolation.
IV Discussion

### v2

@@ Abstract
Background: Residential aged-care facilities (RACFs, also called long-term care facilities, aged care homes, or nursing homes) have elevated risks of respiratory infection outbreaks and associated disease burden. During the COVID-19 pandemic, social isolation policies were commonly used in these facilities to prevent and mitigate outbreaks. We refer specifically to general isolation policies that were intended to reduce contact between residents, without regard to confirmed infection status. Such policies are controversial because of their association with adverse mental and physical health indicators and there is a lack of modelling that assesses their effectiveness.
Methods: In consultation with the Australian Government Department of Health and Aged Care, we developed an agent-based model of COVID-19 transmission in a structured population, intended to represent the salient characteristics of a residential care environment. Using our model, we generated stochastic ensembles of simulated outbreaks and compared summary statistics of outbreaks simulated under different mitigation conditions. Our study focuses on the marginal impact of general isolation (reducing social contact between residents), regardless of confirmed infection. For a realistic assessment, our model included other generic interventions consistent with the Australian Government’s recommendations released during the COVID-19 pandemic: isolation of confirmed resident cases, furlough (mandatory paid leave) of staff members with confirmed infection, and deployment of personal protective equipment (PPE) after outbreak declaration.
Results: In the absence of any asymptomatic screening, general isolation of residents to their rooms reduced median cumulative cases by approximately 27%. However, when conducted concurrently with asymptomatic screening and isolation of confirmed cases, general isolation reduced the median number of cumulative infections by only 12% in our simulations.
Conclusions: Under realistic sets of assumptions, our simulations showed that general isolation of residents did not provide substantial benefits beyond those achieved through screening, isolation of confirmed cases, and deployment of PPE. Our results also highlight the importance of effective case isolation, and indicate that asymptomatic screening of residents and staff may be warranted, especially if importation risk from the outside community is high. Our conclusions are sensitive to assumptions about the proportion of total contacts in a facility accounted for by casual interactions between residents.
I Keywords
residential aged care; nursing homes; long-term care; agent-based model; COVID-19; outbreak response; social isolation; nonpharmaceutical intervention; infectious disease dynamics; contact network
II Background
@@
Residential aged-care facilities (RACFs, also called long-term care facilities, aged care homes, or nursing homes) have elevated risks of respiratory infection outbreaks and associated disease burden. Heightened risk arises due to importation of pathogens by visitors and staff, close and prolonged contact among residents (occupants of the facility) and staff members, and the demographic and health profiles of residents. These intersecting factors produce scenarios in which outbreaks are difficult to mitigate and carry disproportionate consequences in terms of medical impact. Simultaneously, the physical infrastructure, insecure workforce, high health needs of residents, and particular social requirements of RACFs impose limitations on the types of control measures that can be enacted without compromising the mental health of residents [1]. In contrast with other types of healthcare settings, RACFs feature long-term stays (multiple years [2]), frequent social interactions between residents, and a large part-time workforce with variable levels of medical training [3].
@@
Modelling efforts made early during the COVID-19 pandemic did not emphasize the confluence of risk factors present in aged-care environments (e.g., high prevalence of comorbidity and limited infection control capacity) that would later be acknowledged as responsible for large numbers of preventable deaths [4]. This led to a substantial and ongoing global research effort applying infectious disease modelling to outbreak detection, response, and prevention in residential care scenarios [5, 6, 7, 8, 9, 10, 11].
@@
In particular, there is a lack of modelling work that specifically assesses the effectiveness of general isolation conditions applied to resident populations, despite these being possibly the most controversial policies implemented across the sector globally. By “general isolation” we refer specifically to policies that were intended to reduce contact between residents, without regard to confirmed infection status. These include, for example, shutting down shared meal facilities and closing social spaces at facilities, or encouraging residents to stay in their rooms. Around the world, though such policies were not consistently included in government recommendations, many research articles, clinical reports, interviews, media reports, and informal testimony agree that facilities implemented general isolation measures in attempts to reduce the impact of ongoing outbreaks [14, 15, 16, 17, 18, 19].
Here we apply an agent-based model of outbreak mitigation in aged care, to investigate the effectiveness of such policies, and assess the conditions under which they may be justified. For a realistic assessment, our model includes other generic interventions consistent with the Australian Government’s recommendations released during the COVID-19 pandemic: isolation of confirmed resident cases, furlough (mandatory paid leave) of staff with confirmed infection, and deployment of personal protective equipment (PPE) after outbreak declaration [20].
III Methods
III.1 overview
In consultation with the Australian Government Department of Health and Aged Care, we developed an agent-based model of COVID-19 transmission in a structured population, building upon and modifying earlier modelling frameworks used to assess international arrival and quarantine pathways [21]. In general, structured populations feature spatial and temporal clustering in contact patterns between individuals determined by, for example, heterogeneous but persistent dwelling locations and activity schedules. For such populations, homogeneous assumptions about contact patterns are not appropriate for detailed simulations of infectious disease transmission dynamics. To capture these important structural properties, key features of the model include detailed representation of:
@@ III Methods
To capture the limitations of early detection, we implemented realistic screening and response strategies. These were based on those set out in the Australian Government Department of Health document “COVID-19 Outbreaks in Residential Care Facilities, Communicable Disease Network Australia, National Guidelines for the Prevention, Control and Public Health Management of COVID-19 Outbreaks in Residential Care Facilities” dated February 15th, 2022 [20]. Full details of the model implementation and assumptions are provided in the Supporting Information.
III.2 facility structure
Briefly, a facility is modelled as a static population of residents, allocated to single-occupancy rooms. For this study, we simulated a facility with 121 staff and 88 residents, which is a typical size for an Australian RACF [22]. Staff attend a facility according to a weekly roster, with each staff member working five, three, or two days per week which accounts for a total of 73.8 full-time equivalent (FTE) staff (in Australia, 1.0 FTE corresponds to 5 full-time shifts in one week). Staff members are allocated to a set of rooms housing residents they will visit during the days they attend a facility. These allocations are used to generate a network of potential contacts, from which potentially infectious contact events are randomly sampled (Figure 1). In addition to these structured contacts, random contacts between residents in different rooms are included separately to allow for interactions mediated by communal areas and group activities that are not explicitly simulated in the facility model. Infectious contacts are subject to a probabilistic transmission process that depends on the time since the transmitting case was infected and any mitigation measures in place (see Supporting Information).
@@
III.3 within-host model of infection
Infection proceeds through an incubation period, followed by symptom expression and eventual recovery (see the Supporting Information for more details on the model of individual infectiousness trajectories). Each infected individual is potentially infectious from the moment of exposure, through the incubation period and the duration of symptoms. Infectiousness increases until approximately the time of symptom onset, and then declines until recovery. We assume that, on average, 33% of infections never express symptoms and that symptom expression does not alter infectiousness given contact with someone who is susceptible. That is, the time-dependent trajectory of infectiousness is identical for symptomatic and asymptomatic infections. However, case detection following symptom onset may alter contact patterns due to infection control procedures, see below. The assumed asymptomatic fraction of 33% is consistent with global estimates, however, we do not account for heterogeneity by age which could slightly increase the likelihood of case detection in residents relative to staff members. [23, 24, 25, 26].
III.4 baseline transmission dynamics
@@
We calibrated the disease transmission model to produce a reproductive ratio of $R_{0}\approx 2.4$, to simulate a scenario consistent with SARS-CoV-2 transmission in populations with modest immune-derived protection. These conditions correspond approximately to early estimates of the effective reproduction number of the Omicron variant [29]. Further, the reproductive ratio we selected is broadly consistent with the assumptions used in previous modelling studies [8, 30, 5]. See the Supporting Information for more details on calibration.
III.5 initialisation
III.6 outbreak detection and response
The model includes asymptomatic testing of staff and residents, as well as testing on development of symptoms. Infected residents are (imperfectly) isolated from other residents upon detection, but still have contact with staff. Infected staff are furloughed for seven days upon detection, and have no further contact within a facility during the furlough period (Figure 4). As per the Australian Government’s guidelines (dated 15th February, 2022 [20]), an outbreak is declared once two cases have been detected in residents within a five day period or five cases have been detected among staff members over a seven day period (Figure 3). Once an outbreak has been declared, the frequency of testing of staff and residents may increase and infection control measures (e.g., use of PPE) are put in place, reducing the probabilities of transmission among staff and residents (Figure 5). We assume differential effects of these infection control measures depending on the type of contact. For contacts between residents, we assume a relatively limited effectiveness of 20% due to factors such as lower compliance with physical distancing, inconsistent mask wearing, and limited training in the proper use of PPE. For contacts among staff members, we assume a medium effectiveness of 50% due to higher levels of compliance and training with PPE, but inconsistent infection control during contact events not involving residents. For contacts between staff members and residents, we assume the highest level of effectiveness (90%) due to high levels of compliance and training with PPE protocols (including the proper use of N95 masks) for staff members when interacting with residents. These assumptions are broadly in line with literature estimates of PPE efficacy (see, e.g., [31]). See the Supporting Information for more details related to model implementation of outbreak response measures.
Figure 3: Schematic of the transmission process and outbreak declaration after detection of a specified number of cases within a given time window. Outbreaks are declared either from the detection of two cases in residents within five days, or detection of five cases in staff members within seven days.
III.7 simulated policies
We simulated the following set of policies for outbreak detection through asymptomatic screening:
  1. 1.
@@
For each of the above screening policies, we vary the extent to which the resident population (regardless of case status) is socially isolated from one another following declaration of an active outbreak. Here, we present results for 0% (no general isolation), 50% (partial isolation), and 90% (stringent isolation) reductions in background contact rate. Note that residents in case isolation always have their background contact rate reduced by 90% (all scenarios). The assumption that case isolation is imperfect arises from the limitations to isolation reported in residential aged care settings due to the high prevalence of cognitive impairment and impulsive mobility behaviour, as well as the practical and legal limits to non-consensual physical restraint (see e.g., [32, 18, 33]).
@@
This produces a set of 10 scenarios. For each screening strategy (other than the unmitigated outbreak scenario) we perform simulations until 1000 outbreaks have been simulated (the total number of simulations required for this can vary depending on the screening strategy). This produces an outbreak size distribution that is unimodal because it does not include those simulations in which transmission dies out early on and no outbreak is declared. For the unmitigated scenario (in which outbreaks are not declared), a fixed number of independent simulations are performed (10,000 instances). This produces a bimodal distribution of outbreak sizes because it includes those simulations in which transmission dies out early (Figure 6a). We report summary statistics for the second mode (large outbreaks) to facilitate comparison between unmitigated and declared outbreaks. For each scenario, we report the following summary statistics:
@@
  • time to outbreak declaration: the time between introduction of the index case and declaration of the outbreak.
@@
The total number of infections is a generic measure of outbreak severity, but does not account for time-dependent properties of an outbreak. We report the outbreak duration as a measure of severity because costly and burdensome mitigation measures are in place throughout this period. The total number of residents placed into case isolation due to confirmed infection provides an additional measure of severity relevant both to apparent infection numbers and also to detrimental impacts of social isolation. The time to outbreak declaration and the time to first detection do not vary as functions of outbreak response, and provide estimates of the benefits of asymptomatic screening for early detection. The peak staffing deficit (measured in FTE) is a measure of the labour shortage (absenteeism) in the workforce due to furlough of staff. We chose to report the peak FTE deficit rather than the cumulative deficit because it provides a more intuitive measure of the logistical stress experienced within the facility due to personnel shortages. For example, if one staff member working five shifts in one week were furloughed, this would produce 1.0 FTE deficit. If one staff member working two days per week were furloughed, this would produce a 2/5 FTE deficit.
IV Results
IV.1 The effects of asymptomatic screening strategies
@@
IV.2 The effects of general isolation conditions
@@
Our results demonstrate that, while partial and full isolation of the resident population can make a marginal difference in the absence of asymptomatic screening (Figure 8a), it has a negligible effect when combined with asymptomatic screening during outbreaks (Figure 8b). This somewhat surprising finding is due in part to the effectiveness of asymptomatic screening, targeted case isolation, and PPE use, and in part to a fundamental constraint to isolation of aged care residents: that residents must remain in contact with facility staff during outbreaks. Therefore, even if residents are prevented from interacting with one-another, their interactions with staff members can result in transmission. This, combined with the relatively substantial reductions in transmission from the use of PPE and the isolation of confirmed cases, leaves little to be achieved through general isolation of the resident population. For comparison with Figure 8, frequency distributions of cumulative cases for scenarios in which all outbreak mitigation procedures (PPE, case isolation, and asymptomatic testing) are disabled are shown in the Supporting Information Figure S3. This comparison demonstrates that reducing contact between residents can mitigate against transmission and makes a substantial difference to infection totals when other response measures are not implemented.
Because we model general isolation as a reduction in the contact rate between residents, the marginal impact of general isolation is sensitive to the rate of contact between residents relative to the rate of contact involving staff members (which is not affected by general isolation). In our main results, we assume that these proportions are the same (with each rate set to an average of three per resident per day), which is consistent with empirical estimates for residential care facilities (see Discussion). However, our sensitivity analysis demonstrates that if resident-resident contact rates increase, our model produces a corresponding increase in the marginal impact of general isolation measures with lower relative increases observed for lower reproductive ratios (see Supporting Information, Figure S5).
Another potential sensitivity of our model arises from the assumption that compliance with targeted case isolation is fixed at 90%. In our sensitivity analysis (see Supporting Information Figure S6), we demonstrate that our results about the marginal benefits of general isolation are robust to this assumption, as long as the condition is satisfied that general isolation compliance cannot be greater than compliance with targeted case isolation. While outbreak sizes increase when targeted isolation compliance decreases, the above condition prevents general isolation from providing a substantial additional benefit.
@@
screening strategy general isolation level cum. infections outbreak duration peak FTE deficit cum. residents isolated
full asymptomatic screening 90% 14 (11) 11 (8) 3.4 (6.2) 6 (4)
full asymptomatic screening 50% 15 (12) 12 (10) 3.6 (6.2) 6 (5)
full asymptomatic screening 0% 17 (14.5) 14 (13) 3.6 (6) 7 (8)
asymptomatic screening during outbreaks 90% 29 (26) 14 (7) 6.6 (11.8) 12 (11)
asymptomatic screening during outbreaks 50% 31 (28) 15 (10.5) 6.4 (12) 13 (13)
asymptomatic screening during outbreaks 0% 33 (32) 16 (14) 6.6 (12)) 15 (17)
no asymptomatic screening 90% 32.5 (31) 16 (8) 4.4 (6.6) 9 (7)
no asymptomatic screening 50% 37 (37) 20 (12) 4.2 (7.2) 11 (12)
no asymptomatic screening 0% 44.5 (53) 24 (18) 4.4 (6.6) 15 (22)
unmitigated outbreaks - $173^{*}$ (20) - - -
Median over all unmitigated outbreaks which produced more than 80 infections (corresponds to the second mode of the distribution shown in Figure 6a).
Table 1: Summary statistics for model output over a range of mitigation scenarios. Medians values are shown for cumulative infections, outbreak duration, peak FTE deficit, and the cumulative number residents placed into case isolation. Next to each median value, the interquartile range is shown in parenthesis.
V Discussion